# The $\mathbb{S}_{n}$-equivariant Chow polynomial of the Braid matroid

### v2

@@ 5. Numerical applications
We will conclude the paper by showing how to deduce Corollary B from Theorem A. The homomorphism
$\mathrm{rk}:\Lambda\to\mathbb{Q}[\![x]\!],$
taking $p_{1}\to x$ and $p_{n}\to 0$ for $n>1$, satisfies
$\mathrm{rk}(\operatorname{ch}_{n}(V))=\frac{\dim_{\mathbb{Q}}(V)}{n!}.$
Applying this to Theorem A, we obtain
@@
Define the polynomials $\omega_{n}(t)$ inductively by setting $\omega_{1}(t):=1$ and
@@
$\frac{1}{n!}\cdot\frac{\omega_{n}(t-1)}{t-1}.$
This calculation is made via the isomorphism
$\mathcal{M}_{0,n+1}\cong\mathrm{Conf}_{n-2}(\mathbb{P}^{1}\smallsetminus\{0,1,\infty\}).$
Therefore
@@
$t\cdot\mathrm{H}^{\mathrm{num}}_{n}(t)=\sum_{k=1}^{n}\mathrm{Bell}_{n,k}(\omega_{1}(t-1),\ldots,\omega_{n-k+1}(t-1))\cdot\mathrm{H}_{k}^{\mathrm{num}}(t),$
for $n\geq 2$, where $\mathrm{Bell}_{n,k}(x_{1},\ldots,x_{n-k+1})$ is the partial exponential Bell polynomial: see [Com74, §3.3,§3.4]. Subtracting $\mathrm{H}_{n}^{\mathrm{num}}(t)$ from both sides we see that
@@
which is the first part of Corollary B. To obtain the second part of Corollary B, first divide both sides of (5.2) by (t1)(t-1), and then take the limit of both sides as $t\to 1$.
@@
The recursion (5.2) for $\mathrm{H}_{n}^{\mathrm{num}}(t)$ also follows from the recursion in [JKU21, FMSV24]: For a general matroid $M$ with Chow polynomial $\mathrm{H}^{\mathrm{num}}_{M}(t)$ it holds that
@@ 6. Relative stable maps to $\mathbb{P}^{1}$
Translating the definition of $n$-marked combinatorial rubber maps in [Kan23, Definition 2.5] to the present setting, the moduli space $\mathcal{R}_{n}$ is stratified by level trees, just as in the case of $\mathcal{B}_{n}$. Via [Kan23, Equation 2.14] the coarse moduli space corresponding to a level tree factors as a product as in Lemma 2.9. In particular, this leads to an equality
$\mathcal{R}_{n}]=[\mathcal{B}_{n}]$
of the classes of the two moduli spaces in the Grothendieck ring of varieties $K_{0}(\mathsf{Var})$. This equality can also be proved by analyzing the torus fibers of the vertical maps in (6.1). Since both spaces are smooth and proper Deligne–Mumford stacks, it follows that
@@
The equality of dimensions in Theorem 6.3 cannot be upgraded to a space-level isomorphism: the moduli space $\mathcal{R}_{n}$ is only smooth when viewed as a Deligne–Mumford stack, and its coarse moduli space has finite quotient singularities. On the other hand, $\mathcal{B}_{n}$ is a smooth projective variety. We do expect that Theorem 6.3 can be upgraded to an equality of characters of $\mathbb{S}_{n}$-representations, but we do not have a proof.

### v3

@@ 5. Numerical applications
We now show how to deduce Corollary B from Theorem A.
Proof of Corollary B.
The homomorphism
$\mathrm{rk}:\Lambda[\![t]\!]\to\mathbb{Q}[\![x,t]\!],$
taking $p_{1}\to x$ and $p_{n}\to 0$ for $n>1$ satisfies
$\mathrm{rk}(\operatorname{ch}_{n}(V))=\dim_{\mathbb{Q}}(V)\frac{x^{n}}{n!},$
as can be seen directly from the definition of $\operatorname{ch}_{n}(V)$ in (3.4). Moreover, one can verify that if $f,g\in\Lambda[\![t]\!]$ and $g$ is in the ideal $(p_{1},p_{2},\ldots)$, then $\mathrm{rk}(f\circ g)=\mathrm{rk}(f)\circ\mathrm{rk}(g)$, where on the right $\circ$ denotes composition of power series in the variable $x$. Applying this to Theorem A, we obtain
@@
Now define polynomials $\omega_{n}(t)$ inductively by setting $\omega_{1}(t):=1$ and
@@
$\frac{1}{n!}\cdot\frac{\omega_{n}(t-1)}{t-1};$
this equality may be established e.g. by using the isomorphism
$\mathcal{M}_{0,n+1}\cong\mathrm{Conf}_{n-2}(\mathbb{P}^{1}\smallsetminus\{0,1,\infty\})$
and counting the $\mathbb{F}_{q}$-points of $\mathrm{Conf}_{n-2}(\mathbb{P}^{1}\smallsetminus\{0,1,\infty\})$. Therefore
@@
(5.2) $t\cdot\mathrm{H}^{\mathrm{num}}_{n}(t)=\sum_{k=1}^{n}\mathrm{Bell}_{n,k}(\omega_{1}(t-1),\ldots,\omega_{n-k+1}(t-1))\cdot\mathrm{H}_{k}^{\mathrm{num}}(t)$
for $n\geq 2$. Above, $\mathrm{Bell}_{n,k}(x_{1},\ldots,x_{n-k+1})$ is the partial exponential Bell polynomial, defined by the formula
(5.3) $\mathrm{Bell}_{n,k}(x_{1},\ldots,x_{n-k+1})=n!\sum_{\begin{subarray}{c}{\lambda\vdash n}\\ {\ell(\lambda)=k}\end{subarray}}\prod_{i=1}^{n-k+1}\frac{x_{i}^{\lambda_{i}}}{i!^{\lambda_{i}}\lambda_{i}!}$
where for a partition $(1^{\lambda_{1}},2^{\lambda_{2}},\ldots)$ we write $\ell(\lambda)=\sum_{i}\lambda_{i}$; see [Com74, §3.3,§3.4]. Subtracting $\mathrm{H}_{n}^{\mathrm{num}}(t)$ from both sides of (5.2) we see that
@@
which is the first part of Corollary B. To obtain the second part of Corollary B, we divide both sides of (5.4) by (t1)(t-1), and then take the limit of both sides as $t\to 1$. We find
$\frac{\mathrm{Bell}_{n,k}(\omega_{1}(t-1),\ldots,\omega_{n-k+1}(t-1))}{(t-1)}$ $=n!\sum_{\begin{subarray}{c}{\lambda\vdash n}\\ {\ell(\lambda)=k}\end{subarray}}\frac{1}{(t-1)}\prod_{i=1}^{n-k+1}\frac{\omega_{i}(t-1)^{\lambda_{i}}}{i!^{\lambda_{i}}\lambda_{i}!}$
$=n!\sum_{\begin{subarray}{c}{\lambda\vdash n}\\ {\ell(\lambda)=k}\end{subarray}}\frac{1}{\lambda_{1}!}\cdot\frac{1}{(t-1)}\prod_{i=2}^{n-k+1}\frac{(t-1)^{\lambda_{i}}\cdots(t-i+1)^{\lambda_{i}}}{i!^{\lambda_{i}}\lambda_{i}!}.$
Direct inspection of the above expression shows that the summand corresponding to $\lambda\vdash n$ has nonzero limit as $t\to 1$ if and only if $\lambda=(1^{\lambda_{1}},j^{1})$ for some $j$. Moreover, since $\ell(\lambda)=k$, we must have $\lambda_{1}+1=k$, so $\lambda_{1}=k-1$, and j=nk+1j=n-k+1. Therefore we conclude that
$\lim_{t\to 1}\frac{\mathrm{Bell}_{n,k}(\omega_{1}(t-1),\ldots,\omega_{n-k+1}(t-1))}{(t-1)}=n!\cdot\frac{1}{(k-1)!}\frac{(-1)^{n-k-1}(n-k-1)!)}{(n-k+1)!}.$
This expression, together with (5.4), implies the second part of Corollary B. ∎
@@
The recursion (5.4) for $\mathrm{H}_{n}^{\mathrm{num}}(t)$ also follows from the recursion in [JKU21, FMSV24]: For a general matroid $M$ with Chow polynomial $\mathrm{H}^{\mathrm{num}}_{M}(t)$ we have
@@ 6. Relative stable maps to $\mathbb{P}^{1}$
Translating the definition of $n$-marked combinatorial rubber maps in [Kan23, Definition 2.5] to the present setting, the moduli space $\mathcal{R}_{n}$ is stratified by level trees, just as in the case of $\mathcal{B}_{n}$. Via [Kan23, Equation 2.14] the coarse moduli space corresponding to a level tree factors as a product as in Lemma 2.12. In particular, this leads to an equality $\mathcal{R}_{n}]=[\mathcal{B}_{n}]$ of the classes of the two moduli spaces in the Grothendieck ring of varieties $K_{0}(\mathsf{Var})$. This equality can also be proved by analyzing the torus fibers of the vertical maps in (6.1). Since both spaces are smooth and proper Deligne–Mumford stacks, it follows that
@@
By [Kan23, Theorem B], Theorem 6.3 also holds if $\mathcal{R}_{n}$ is replaced by any moduli space of relative stable maps $\overline{\mathcal{M}}(\boldsymbol{x})$ where $\boldsymbol{x}\in(\mathbb{Z}\smallsetminus\{0\})^{n+1}$ has exactly one positive entry.
Remark 6.5.
The equality of dimensions in Theorem 6.3 cannot be upgraded to a space-level isomorphism: the moduli space $\mathcal{R}_{n}$ is only smooth when viewed as a Deligne–Mumford stack, and its coarse moduli space has finite quotient singularities. On the other hand, $\mathcal{B}_{n}$ is a smooth projective variety. We do expect that Theorem 6.3 can be upgraded to an equality of characters of $\mathbb{S}_{n}$-representations. It is also natural to ask whether the equality of Chow groups in Theorem 6.3 can be upgraded to an isomorphism of Chow rings. We expect that the two rings are not isomorphic.